# Convergence of the open WASEP stationary measure without Liggett’s condition

### v1

@@ 2.4. Notation for Asymptotics
We will use the following notation to simplify the expressions for the logarithms of the $q$-pochammer symbols.
(2.13) $\mathcal{A}^{+}[\kappa,z]=-\frac{\pi^{2}}{6\kappa}-\left(z-\frac{1}{2}\right)\log{(\kappa)}-\log{\left[\frac{\Gamma(z)}{\sqrt{2\pi}}\right]}$ $\mathcal{A}^{-}[\kappa,z]=\frac{\pi^{2}}{12\kappa}-\left(z-\frac{1}{2}\right)\log{(2)}.$
@@
$\log{\left[(q^{z};q)_{\infty}\right]}=\mathcal{A}^{+}[\kappa,z]-\sum_{n=1}^{m-1}\frac{B_{n+1}(z)B_{n}}{n(n+1)!}\kappa^{n}+\textsc{Error}^{+}_{m}[\kappa,z],$
$\log{\left[(-q^{z};q)_{\infty}\right]}=\mathcal{A}^{-}[\kappa,z]-\sum_{n=1}^{m-1}(2^{n}-1)\frac{B_{n+1}(z)B_{n}}{n(n+1)!}\kappa^{n}+\textsc{Error}^{-}_{m}[\kappa,z],$
where $B_{n}(z)$ and $B_{n}$ denote Bernoulli polynomials and Bernoulli numbers, respectively, and $\textsc{Error}_{m}^{\pm}[\kappa,z]$ are error terms. For any $\alpha\in(0,\pi)$, $\epsilon\in(0,1/2)$ and any $b\in(m-1,m)$ there exist $C,\kappa_{0}>0$ such that for all $\kappa\in(0,\kappa_{0})$ and all $z\in\mathbb{C}$ with $|Im(z)|<\frac{\alpha}{\kappa}$,
(2.14) $|\textsc{Error}_{1}^{\pm}[\kappa,z]|\leq C(\kappa(1+|z|)^{2}+\kappa^{b}(1+|z|)^{1+2b+\epsilon}).$
[…]
@@
$\mathcal{A}^{\pm}_{N}[z_{1},...,z_{k}]=\sum_{i=1}^{k}\mathcal{A}^{\pm}\left[\kappa,z_{i}\right],$ $E^{\pm}_{N}[z_{1},...,z_{k}]=\sum_{i=1}^{k}\textsc{Error}_{1}^{\pm}\left[\kappa,z_{i}\right].$
@@ 2.5. Convergence of the Laplace transform
For a collection of random vectors of length $d$, $\{\mathbf{X}^{(i)}\}_{i\in\mathbb{N}}$, with $\mathbf{X}^{(i)}:=(X^{(i)}_{1},\ldots,X^{(i)}_{d})$ and $\mathbf{z}\in\mathbb{R}^{d}$ defined so that $\mathbf{z}:=(z_{1},\ldots,z_{d})$, we define the collection of functions $\{L_{i}(\mathbf{z})\}_{i\in\mathbb{N}}$ to be the Laplace transforms of each of the $\mathbf{X}^{(i)}$, respectively,
$L_{i}(\textbf{z})=\mathbb{E}\left[\exp{\left(\sum_{j=1}^{d}z_{j}X_{j}^{(i)}\right)}\right].$
[…]
3. Convergence of the Askey-Wilson Densities
The purpose of this section and the next is to demonstrate that the sequence of Laplace transforms of the Askey-Wilson process, under the scaling assumptions without Liggett’s condition, given in 1.1, converges pointwise to the Laplace transform of the open KPZ stationary measure.
@@ 3. Convergence of the Askey-Wilson Densities
[…]
(3.1) $\varphi^{(N)}_{u,v,w,r}(\textbf{c},\textbf{x})$ $=\frac{\mathbb{E}\left[\prod_{k=1}^{d+1}\left(\cosh(s_{k}^{(N)})+\mathbb{Y}_{e^{-2s_{k}^{(N)}}}\right)^{n_{k}-n_{k-1}}\right]}{\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]}.$
For all $\textbf{c}\in(0,C_{u,r})^{d}$, defined so that
$C_{u,r}$ :=$:=\begin{cases}2&u\leq 0,r\leq 0\\ 2r&r>0,u\leq 0\\ \min{\{2r,2u\}}&u,r>0\end{cases}$
, and $u+v>0$ and $w,r>0$, this sequence converges point-wise to a function $\varphi_{u,v}(\textbf{c},\textbf{x})$, defined
(3.2) $\varphi_{u,v}(\textbf{c},\textbf{x})$ $:=\frac{\mathbb{E}\left[\exp{\left(\frac{1}{4}\sum_{k=1}^{d+1}(s_{k}^{2}-\mathbb{T}_{s_{k}})(x_{k}-x_{k-1})\right)}\right]}{\mathbb{E}\left[\exp{\left(-\frac{1}{4}\mathbb{T}_{0}\right)}\right]},$
where $\mathbb{T}_{t}$ is the continuous dual Hahn process. Moreover, $\varphi_{u,v}(\textbf{c},\textbf{x})$ is the Laplace transform of the stationary measure of the open KPZ equation constructed in [CK24].
The first step in the proof of 3.1 is to demonstrate the convergence of the Askey-Wilson process measures to those of the continuous dual Hahn process, which we do in the rest of this section. In Section 4 we will complete the proof of 3.1. Finally, in Section 5, we give the proof of 1.3.
We will use the notation $x_{j}^{\chi}(t)$ and $y_{j}^{(N),\chi}$, respectively, for the locations of the atoms of the continuous dual Hahn process and Askey-Wilson process (at the $N$-th step of scaling) associated to $\chi\in\{q^{v+t/2},q^{u-t/2},-q^{w+t/2},-q^{r-t/2}\}$.
@@
Under the scaling of the Askey-Wilson measure when $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)}$, see 2.1,
@@
    For all $y\in\hat{S}^{c}$ and $t\in\mathbb{R}_{>0}$, for all $\eta>1$ there exist $N_{0}\in\mathbb{N}$ and $C,\chi>0$ so that for all $N>N_{0}$ and $y\in[0,4N]$,
    $N^{u+v}\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t}^{(N),c}(y)=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{c}(y)\cdot e^{\textsc{Error}_{t}^{(N),c}(y)},$
@@
    $\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{(N),c}(y)$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{c}(y),$
@@
    Similarly, for $\hat{y}_{j}^{(N),\chi}(t)\in\hat{S}^{d}_{t}$ and $t\in(-2w,2r)$, there exist $C_{1},C_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{>0}$ such that for all $N\in\mathbb{Z}_{\geq 1}$, the atoms $y_{j}^{(N),\chi}$ for $\chi\in\{q^{v+\frac{t}{2}},q^{u-\frac{t}{2}}\}$ satisfy
    $N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}^{(N),v}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))\cdot e^{\textsc{Error}_{t}^{(N),d}(x_{j}^{v}(t))},$
    $N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}^{(N),u}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{u}(t))\cdot e^{\textsc{Error}_{t}^{(N),d}(x_{j}^{u}(t))},$
@@
    $\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),\chi}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{\chi}(t)).$
The limiting measure $\mathfrak{p}_{t}^{c}(y)$ and $\mathfrak{p}_{t}^{d}(y)$ are the one-point measures of the continuous dual Hahn process, see Section 2.3.
Similarly, we can compute the limit of $\hat{\pi}_{s,t}^{(N),c,c}(\cdot,\cdot)$ and $\hat{\pi}_{s,t}^{(N),d,c}(\cdot,\cdot)$ under scaling.
@@
With scaling conditions as in 1.1, and for all $s<t$
@@
    For all $m,r\in\hat{S}^{(N),c}=[0,4N]$ and $\eta>0$, there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi\in\mathbb{R}_{>0}$ such that for all $N>N_{0}$
    $\sqrt{1-\frac{r}{4N}}\hat{\pi}_{s,t}^{(N),c,c}(m,r)=\mathfrak{p}_{s,t}^{c,c}(m,r)\cdot e^{\textsc{Error}_{s,t}^{(N),c,c}(m,r)}$
@@
    $|\textsc{Error}_{s,t}^{(N),c,c}(m,r)|\leq CN^{-\chi}(1+\sqrt{m})^{\eta}+CN^{-\chi}(1+\sqrt{r})^{\eta}.$
@@
    If $s,t\in(-2b,2d)$ and $v+\frac{s}{2}<0$ so that $\hat{S}_{s}^{(N),d}$ is entirely composed of $v$ atoms $\hat{y}_{j}^{(N),v}(s)$ for $j\in[[0,\lfloor-v-\frac{s}{2}\rfloor]]$ then for each such $j$, and all $r\in\hat{S}^{(N),c}=[0,4N]$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi\in\mathbb{R}_{>0}$ such that for all $N>N_{0}$
    $\hat{\pi}_{s,t}^{(N),d,c}\left(\hat{y}_{j}^{(N),v}(s),r\right)=\mathfrak{p}_{s,t}^{d,c}(x_{j}^{v}(s),r)\cdot e^{\textsc{Error}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)}$
@@
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,c}(\hat{y}_{j}^{(N),v}(s),r)=\mathfrak{p}_{s,t}^{d,c}(x_{k}^{v}(s),r).$
@@
    If $s,t\in(-2w,2r)$ and $u-\frac{s}{2}<0$ so that $\hat{S}_{s}^{(N),d}$ is entirely composed of $u$ atoms $\hat{y}_{j}^{(N),u}(s)$ for $j\in[[0,\lfloor-u+\frac{s}{2}\rfloor]]$ then for each such $j$ and for all $r\in\hat{S}^{(N),c}=[0,4N]$,
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,c}(\hat{y}_{j}^{(N),v}(s),r)=0$
Finally, we obtain the following asymptotic result for $\hat{\pi}_{t}^{(N),d,d}(\cdot,\cdot)$ and $\hat{\pi}_{t}^{(N),c,d}(\cdot,\cdot)$.
@@
    If $v+\frac{s}{2}<0$ then the discrete support of $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\cdot)$ coincides with the points $\hat{y}_{k}^{(N),v}(t)$ for $k\in[[0,\lfloor-v-\frac{t}{2}\rfloor]]$ and $\hat{y}_{k}^{(N),u}(t)$ for $k\in[[0,\lfloor-u+\frac{t}{2}\rfloor]]$ and there exist $C,\chi>0$ such that for all $N>0$ and $k\in[[0,\lfloor-v-\frac{t}{2}\rfloor]],$ respectively, $k\in[[0,\lfloor-u+\frac{t}{2}\rfloor]]$
    $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\hat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{v}(t))\cdot e^{\textsc{Error}_{s,t}^{(N),d,d}(v,v)}$
    $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\hat{y}_{k}^{(N),u}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{u}(t))\cdot e^{\textsc{Error}_{s,t}^{(N),d,d}(v,u)}$
    and the error terms are bounded so that
@@
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\hat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{v}(t)).$
@@
    If $u-\frac{s}{2}<0$ then the discrete support of $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),u}(s),\cdot)$ does not contain any atoms.
@@
    If $u-\frac{s}{2}<0$ then the discrete support of $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),u}(s),\cdot)$ coincides with the points $\hat{y}_{k}^{(N),v}(t)$ for $k\in[[0,\lfloor-v-\frac{t}{2}\rfloor]]$ and
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),u}(s),\hat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{u}(s),x_{k}^{v}(t))$
  4. (4)
    If $v+\frac{s}{2}<0$ and $r\in\hat{S}^{(N),c}=[0,4N]$ then the Askey-Wilson measure $\hat{\pi}_{s,t}^{(N),c,d}(r,\cdot)$ has no atomic part. When $u-\frac{s}{2}<0$ and $r\in\hat{S}^{(N),c}=[0,4N]$, the measure $\hat{\pi}_{s,t}^{(N),c,d}(r,\cdot)$ has discrete support on the atoms $\hat{y}_{k}^{(N),u}(t)$ and there exist $C,\chi\in\mathbb{R}_{>0}$ such that for all $N\in\mathbb{Z}_{\geq 1}$,
    $\hat{\pi}_{s,t}^{(N),c,d}\left(r,\hat{y}_{k}^{(N),u}(t)\right)=\mathfrak{p}_{s,t}^{c,d}(r,x_{k}^{u}(t))\cdot e^{\textsc{Error}_{s,t}^{c,d}(r,x_{k}^{u}(t))}$
@@
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),c,d}(r,\hat{y}_{k}^{(N),u}(t))=\mathfrak{p}_{s,t}^{c,d}(r,x_{k}^{u}(t)).$
3.1. Convergence of One-point Distributions

### v2

@@ 2.4. Notation for Asymptotics
We will use the following notation to simplify the expressions for the logarithms of the $q$-Pochhammer symbols.
$\mathcal{A}^{+}(\kappa,z)=-\frac{\pi^{2}}{6\kappa}-\left(z-\frac{1}{2}\right)\log{(\kappa)}-\log{\left(\frac{\Gamma(z)}{\sqrt{2\pi}}\right)},$ $\mathcal{A}^{-}(\kappa,z)=\frac{\pi^{2}}{12\kappa}-\left(z-\frac{1}{2}\right)\log{(2)}.$ (2.13)
@@
$\log{\left((q^{z};q)_{\infty}\right)}=\mathcal{A}^{+}(\kappa,z)-\sum_{n=1}^{m-1}\frac{B_{n+1}(z)B_{n}}{n(n+1)!}\kappa^{n}+e^{+}_{m}(\kappa,z),$
$\log{\left((-q^{z};q)_{\infty}\right)}=\mathcal{A}^{-}(\kappa,z)-\sum_{n=1}^{m-1}(2^{n}-1)\frac{B_{n+1}(z)B_{n}}{n(n+1)!}\kappa^{n}+e^{-}_{m}(\kappa,z),$
where $B_{n}(z)$ and $B_{n}$ denote Bernoulli polynomials and Bernoulli numbers, respectively. For any $\alpha\in(0,\pi)$, $\epsilon\in(0,1/2)$ and any $b\in(m-1,m)$ there exist $C,\kappa_{0}>0$ such that for all $\kappa\in(0,\kappa_{0})$ and all $z\in\mathbb{C}$ with $|Im(z)|<\frac{\alpha}{\kappa}$,
$|e_{1}^{\pm}(\kappa,z)|\leq C(\kappa(1+|z|)^{2}+\kappa^{b}(1+|z|)^{1+2b+\epsilon}).$ (2.14)
The bound on $e^{+}_{m}(\kappa,z)$ further holds with the condition $|Im(z)|<\frac{\alpha}{\kappa}$ replaced by $|Im(z)|<\frac{2\alpha}{\kappa}$. Furthermore, for any $r>0$, $\epsilon\in(0,1/2)$ and $b\in(m-1,m)$ there exist $C,\kappa_{0}>0$ such that for all $\kappa\in(0,\kappa_{0})$ and all $z\in\mathbb{C}$ with dist$(\text{Re}(z),\mathbb{Z}_{\leq 0})>r$, (2.14) continues to hold.
@@
$\mathcal{A}^{\pm}_{N}\left(z_{1},...,z_{k}\right)=\sum_{i=1}^{k}\mathcal{A}^{\pm}\left(\kappa,z_{i}\right),$ $E^{\pm}_{N}\left(z_{1},...,z_{k}\right)=\sum_{i=1}^{k}e_{1}^{\pm}\left(\kappa,z_{i}\right).$ (2.15)
@@ 2.5. Convergence of the Laplace transform
For a collection of random vectors of length $d$, $\{\vec{X}^{(i)}\}_{i\in\mathbb{N}}$, with $\vec{X}^{(i)}:=(X^{(i)}_{1},\ldots,X^{(i)}_{d})$ and $\vec{z}\in\mathbb{R}^{d}$ defined so that $\vec{z}:=(z_{1},\ldots,z_{d})$, we define the collection of functions $\{L_{i}(\vec{z})\}_{i\in\mathbb{N}}$ to be the Laplace transforms of each of the $\vec{X}^{(i)}$, respectively,
$L_{i}(\vec{z})=\mathbb{E}\left[\exp{\left(\sum_{j=1}^{d}z_{j}X_{j}^{(i)}\right)}\right].$
Suppose that $L_{i}(\vec{z})$ are all finite and converge pointwise to a function $L(\vec{z})$ for all $\vec{z}$ in an open set of $\mathbb{R}^{d}$. If $L(\vec{z})$ is the Laplace transform of a random variable, $\vec{Y}=(Y_{1},...,Y_{d})$, then the sequence $\vec{X}^{(i)}$ converges in distribution to $\vec{Y}.$
3. Convergence of the Askey Wilson Tangent Laws to the Continuous Dual Hahn Laws
The purpose of this section and the next is to demonstrate that the sequence of Laplace transforms of the Askey-Wilson process, under 1.1, converges pointwise to the Laplace transform of the open KPZ stationary measure.
@@ 3. Convergence of the Askey Wilson Tangent Laws to the Continuous Dual Hahn Laws
The multi-point Laplace transforms $\varphi^{(N)}_{u,v,w,r}(\vec{c},\vec{x})$, with $\vec{c},\vec{x}$ as in (2.2), of the sequence of open ASEP height functions $\{h_{u,v,w,r}^{(N)}(\cdot)\}_{N\in\mathbb{N}}$, scaled as in 1.1, are represented in terms of the Askey-Wilson process $\mathbb{Y}_{q^{s}}$ as
$\varphi^{(N)}_{u,v,w,r}(\vec{c},\vec{x})$ $=\frac{\mathbb{E}\left[\prod_{k=1}^{d+1}\left(\cosh(s_{k}^{(N)})+\mathbb{Y}_{e^{-2s_{k}^{(N)}}}\right)^{n_{k}-n_{k-1}}\right]}{\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]}.$ (3.1)
For all $\vec{c}\in(0,C_{u,r})^{d}$, defined so that
$C_{u,r}$ :=$:=\begin{cases}2&u\leq 0,r\leq 0,\\ 2r&r>0,u\leq 0,\\ \min{\{2r,2u\}}&u,r>0,\end{cases}$
for $u+v>0$ and $w,r>0$, we have pointwise convergence $\lim_{N\to\infty}\phi^{(N)}_{u,v,w,r}(\vec{c},\vec{x})=\varphi_{u,v}(\vec{c},\vec{x})$. The function $\varphi_{u,v}(\vec{c},\vec{x})$ is defined
$\varphi_{u,v}(\vec{c},\vec{x})$ $:=\frac{\mathbb{E}\left[\exp{\left(\frac{1}{4}\sum_{k=1}^{d+1}(s_{k}^{2}-\mathbb{T}_{s_{k}})(x_{k}-x_{k-1})\right)}\right]}{\mathbb{E}\left[\exp{\left(-\frac{1}{4}\mathbb{T}_{0}\right)}\right]},$ (3.2)
where $\mathbb{T}_{t}$ is the continuous dual Hahn process. Moreover, $\varphi_{u,v}(\vec{c},\vec{x})$ is the Laplace transform of the stationary measure of the open KPZ equation constructed in [CK24].
The first step in the proof of 3.1 is to demonstrate the convergence of the Askey-Wilson process measures to those of the continuous dual Hahn process, which we do in the rest of this section. In Section 4 we will complete the proof of 3.1. Finally, in Section 5, we give the proof of 1.1.
We will use the notation $x_{j}^{\theta}(t)$ and $y_{j}^{(N),\theta}$, respectively, for the locations of the atoms of the rescaled continuous dual Hahn process and Askey-Wilson process associated to $\theta\in\{v,u,w,r\}$ (with correspond to the atoms arising from, respectively, $\{q^{v+t/2},q^{u-t/2},-q^{w+t/2},-q^{r-t/2}\}$).
@@
Under the scaling of the Askey-Wilson measure when $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)}$, the following convergence results hold.
@@
    For all $y\in\widehat{S}^{c}$ and $t\in\mathbb{R}_{>0}$, for all $\eta>1$ there exist $N_{0}\in\mathbb{N}$ and $C,\chi>0$ so that for all $N>N_{0}$ and $y\in[0,4N]$,
    $N^{u+v}\sqrt{1-\frac{y}{4N}}\widehat{\pi}_{t}^{(N),c}(y)=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{c}(y)\cdot e^{\textsc{Error}_{t}^{(N),c}(y)},$
@@
    $\lim_{N\to\infty}N^{u+v}\widehat{\pi}_{t}^{(N),c}(y)$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{c}(y),$
@@
    Similarly, for $\widehat{y}_{j}^{(N),\theta}(t)\in\widehat{S}^{d}_{t}$ and $t\in(-2w,2r)$, there exist $C_{1},C_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{>0}$ such that for all $N\in\mathbb{Z}_{\geq 1}$, the atoms $y_{j}^{(N),\theta}(t)$ for $\theta\in\{v,u\}$ satisfy
    $N^{u+v}\widehat{\pi}_{t}^{(N),d}(\widehat{y}^{(N),v}_{j}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))\cdot e^{\textsc{Error}_{t}^{(N),d}(x_{j}^{v}(t))},$
    $N^{u+v}\widehat{\pi}_{t}^{(N),d}(\widehat{y}^{(N),u}_{j}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{u}(t))\cdot e^{\textsc{Error}_{t}^{(N),d}(x_{j}^{u}(t))},$
@@
    $\lim_{N\to\infty}N^{u+v}\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{j}^{(N),\chi}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{\chi}(t)).$
The limiting measure $\mathfrak{p}_{t}^{c}(y)$ and $\mathfrak{p}_{t}^{d}(y)$ are the marginal measures of the continuous dual Hahn process, see Section 2.3.
Similarly, we can compute the limit of $\widehat{\pi}_{s,t}^{(N),c,c}(\cdot,\cdot)$ and $\widehat{\pi}_{s,t}^{(N),d,c}(\cdot,\cdot)$ under scaling.
@@
With scaling conditions as in 1.1, and for all $s<t$, the following results hold.
@@
    For all $b,r\in\widehat{S}^{(N),c}=[0,4N]$ and $\eta>0$, there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi\in\mathbb{R}_{>0}$ such that for all $N>N_{0}$
    $\sqrt{1-\frac{r}{4N}}\widehat{\pi}_{s,t}^{(N),c,c}(b,r)=\mathfrak{p}_{s,t}^{c,c}(b,r)\cdot e^{\textsc{Error}_{s,t}^{(N),c,c}(b,r)},$
@@
    $|\textsc{Error}_{s,t}^{(N),c,c}(b,r)|\leq CN^{-\chi}(1+\sqrt{b})^{\eta}+CN^{-\chi}(1+\sqrt{r})^{\eta}.$
@@
    If $s,t\in(-2b,2d)$ and $v+\frac{s}{2}<0$, so that $\widehat{S}_{s}^{(N),d}=\{\widehat{y}_{j}^{(N),v}(s)\}_{j\in[[0,\lfloor-v-\frac{s}{2}\rfloor]]}$, then for each such $j$, and all $r\in\widehat{S}^{(N),c}=[0,4N]$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi\in\mathbb{R}_{>0}$ such that for all $N>N_{0}$
    $\widehat{\pi}_{s,t}^{(N),d,c}\left(\widehat{y}_{j}^{(N),v}(s),r\right)=\mathfrak{p}_{s,t}^{d,c}(x_{j}^{v}(s),r)\cdot e^{\textsc{Error}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)},$
@@
    $\lim_{N\to\infty}\widehat{\pi}_{s,t}^{(N),d,c}(\widehat{y}_{j}^{(N),v}(s),r)=\mathfrak{p}_{s,t}^{d,c}(x_{k}^{v}(s),r).$
@@
    If $s,t\in(-2w,2r)$ and $u-\frac{s}{2}<0$, so that $\widehat{S}_{s}^{(N),d}=\{\widehat{y}_{j}^{(N),u}(s)\}_{j\in[[0,\lfloor-u+\frac{s}{2}\rfloor]]}$, then for each such $j$ and for all $r\in\widehat{S}^{(N),c}=[0,4N]$,
    $\lim_{N\to\infty}\widehat{\pi}_{s,t}^{(N),d,c}(\widehat{y}_{j}^{(N),v}(s),r)=0.$
Finally, we obtain the following asymptotic result for $\widehat{\pi}_{t}^{(N),d,d}(\cdot,\cdot)$ and $\widehat{\pi}_{t}^{(N),c,d}(\cdot,\cdot)$.
@@
    If $v+\frac{s}{2}<0$ then the discrete support of $\widehat{\pi}_{s,t}^{(N),d,d}(\widehat{y}_{j}^{(N),v}(s),\cdot)$ coincides with the points $\widehat{y}_{k}^{(N),v}(t)$ for $k\in[[0,\lfloor-v-\frac{t}{2}\rfloor]]$ and there exist $C,\chi>0$ such that for all $N>0$ and $j\in[[0,\lfloor-v-\frac{s}{2}\rfloor]],$ respectively, $k\in[[0,\lfloor-u+\frac{t}{2}\rfloor]]$
    $\widehat{\pi}_{s,t}^{(N),d,d}(\widehat{y}_{j}^{(N),v}(s),\widehat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{v}(t))\cdot e^{\textsc{Error}_{s,t}^{(N),d,d}(v,v)},$
    and
@@
    $\lim_{N\to\infty}\widehat{\pi}_{s,t}^{(N),d,d}(\widehat{y}_{j}^{(N),v}(s),\widehat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{v}(t)).$
@@
    If $u-\frac{s}{2}<0$ then the discrete support of $\widehat{\pi}_{s,t}^{(N),d,d}(\widehat{y}_{j}^{(N),u}(s),\cdot)$ coincides with the points $\widehat{y}_{k}^{(N),u}(t)$ for $k\in[[0,j]]$ and for all such $j,k$ and the associated atoms
    $\widehat{\pi}_{s,t}^{(N),d,d}\left(\widehat{y}_{j}^{(N),u}(s),\widehat{y}_{k}^{(N),u}(t)\right)$ $=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{u}(s),x_{j}^{v}(t))\cdot e^{\textsc{Error}_{s,t}^{(N),d,d}(u,u)},$
    and
    $|\textsc{Error}_{s,t}^{(N),d,d}(\cdot,\cdot)|$ $\leq CN^{-\chi}.$
    In particular,
    $\lim_{N\to\infty}\widehat{\pi}_{s,t}^{(N),d,d}(\widehat{y}_{j}^{(N),u}(s),\widehat{y}_{k}^{(N),u}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{u}(s),x_{k}^{u}(t)).$
@@
    If $v+\frac{s}{2}<0$ and $r\in\widehat{S}^{(N),c}=[0,4N]$ then the measure $\widehat{\pi}_{s,t}^{(N),c,d}(r,\cdot)$ has no atomic part. When $u-\frac{s}{2}<0$ and $r\in\widehat{S}^{(N),c}=[0,4N]$, the measure $\widehat{\pi}_{s,t}^{(N),c,d}(r,\cdot)$ has discrete support on the atoms $\widehat{y}_{k}^{(N),u}(t)$ for $k\in[[0,\lfloor-u+\frac{t}{2}\rfloor]]$ and there exist $C,\chi\in\mathbb{R}_{>0}$ such that for all $N\in\mathbb{Z}_{\geq 1}$,
    $\widehat{\pi}_{s,t}^{(N),c,d}\left(r,\widehat{y}_{k}^{(N),u}(t)\right)=\mathfrak{p}_{s,t}^{c,d}(r,x_{k}^{u}(t))\cdot e^{\textsc{Error}_{s,t}^{c,d}(r,x_{k}^{u}(t))},$
@@
    $\lim_{N\to\infty}\widehat{\pi}_{s,t}^{(N),c,d}(r,\widehat{y}_{k}^{(N),u}(t))=\mathfrak{p}_{s,t}^{c,d}(r,x_{k}^{u}(t)).$
3.1. Convergence of Marginal Distributions